a bullet is fired from a rifle with a muzzle velocity of 80 meters per second at an angle of 30 degrees from the horizontal. How far away does the bullet hit the ground? How long does it take the bullet to hit the ground? What is the maximum height of the bullet's path?
Answer #1

horizontal distance travelled by the bullet

R = U ^2 sin 2 (theta ) / g

= ( 80 m/s ) ^2 ( sin 60) / 9.8 = 565.56 m

time

t = 2 u sin (30 ) / g = 2 (80 m/s ) (0.5 ) / 9.8 = 8.16 s

height

H = u^2 sin 30 ^ 2 / 2 g = ( 80 m/s ) ( sin 30^ 2 / 2 ( 9.8 ) = 81.63 m